ON THE MUMFORD-TATE CONJECTURE FOR ABELIAN

VARIETIES WITH REDUCTI ON CONDITIONS

Alex Lesin

Abstract. We study monodromy action on abelian varieties satisfying certain bad reduction conditions. These conditions allow us to get some control over the Galois image. As a consequence we verify the Mumford–Tate conjecture for such abelian varieties.

Typeset by AMS-TEX